Please help. (:
let f(x)=1-x^2 , g(x) 1-x. Find (f-g)(x)

Answer by stanbon(75887)

Please help. (:
let f(x)=1-x^2 , g(x) 1-x. Find (f-g)(x)
---------
(f-g)(x) = f(x)-g(x)
---
= (1-x^2)-(1-x)
----
= x - x^2
-------------------
Cheers,
Stan H.